### v1

# Semidefinite programming bounds on fractional cut-cover and maximum 2-SAT for highly regular graphs

Abstract
We use semidefinite programming to bound the fractional cut-cover parameter of graphs in association schemes in terms of their smallest eigenvalue. We also extend the equality cases of a primal-dual inequality involving the Goemans-Williamson semidefinite program, which approximates maxcut, to graphs in certain coherent configurations. Moreover, we obtain spectral bounds for max 2-sat when the underlying graphs belong to a symmetric association scheme by means of a certain semidefinite program used to approximate quadratic programs, and we further develop this technique in order to explicitly compute the optimum value of its gauge dual in the case of distance-regular graphs.
1 Introduction
Let $G=(V,E)$ be a simple, undirected, loopless graph with adjacency matrix $A$. The well-known maximum cut problem (maxcut) consists of finding a partition of $V$ into two sets that maximizes the number of edges between them, and we can express this problem as the following quadratic program:
$\text{mc}(G):=\max\left\{\frac{1}{4}\langle L,xx^{T}\rangle:x\in\mathbb{R}^{V},x_{i}^{2}=1\right\},$ (1)
[…]
$\begin{split}\eta(G):=&\max\left\{\frac{1}{4}\langle L,M\rangle:M\succcurlyeq 0,\text{diag}(M)=\mathbbm{1}\right\}\\ =&\min\left\{\rho:\rho\geq 0,x\in\mathbb{R}^{V},\rho\geq\mathbbm{1}^{T}x,\text{Diag}(x)\succcurlyeq\frac{1}{4}L\right\},\end{split}$ (2)
where $X\succcurlyeq Y$ denotes that the square matrix $X-Y$ is positive semidefinite, $\mathbbm{1}$ is the all-ones vector, diag is the function that maps the diagonal entries of a matrix to a vector, Diag is its adjoint that maps a vector to a diagonal matrix containing its entries in its diagonal. The second equality follows from semidefinite programming strong duality.
Much information about $\text{mc}(G)$ and $\eta(G)$ can be obtained by studying other programs that satisfy a type of duality known as antiblocking duality, or more generally as gauge duality. The former has been successfully exploited to study many combinatorial objects throughout the past decades, such as matroids and permutations [10], and the latter has been recently employed as a general framework for studying pairs of graphs parameters such as the coclique number, the fractional chromatic number, and the Lovász Theta [14].
In order to properly define the notion of antiblocking or gauge duality to $\text{mc}(G)$ or $\eta(G)$ it is necessary to introduce them as functions whose input is the graph together with edge weights (see [1] for the specific treatment, or [15, Part III] for a general presentation). In this work, however, we focus on the unweighted scenario (all weights equal to $1$), and therefore we take the shortcut of defining the dual graph parameters as the special case of the dual gauge functions taken on weights everywhere equal to $1$.
The antiblocking dual parameter to maxcut is the fractional cut-cover number (fcc), which is defined as
$\text{fcc}(G):=\min\left\{\mathbbm{1}^{T}y:y\in\mathbb{R}^{\mathcal{P}(V)}_{+},\sum_{S\subseteq V}y_{S}\cdot\mathbbm{1}_{\delta(S)}\geq\mathbbm{1}\right\},$ (3)
where $\mathcal{P}(V)$ is the power set of $V$, $\delta(S)$ is the set of edges with one end in $S$ and the other in $V\setminus S$, and $\mathbbm{1}_{\delta(S)}$ is the indicator vector of $\delta(S)$. Intuitively, the integer solutions of this program will give a cover of the edges of the graph by cuts such that each edge $e$ is covered at least once. In [1], the authors show how to construct an efficient approximation algorithm for $\text{fcc}(G)$ with factor $1/\alpha_{\text{GW}}$, by combining the Goemans-Williamson algorithm with the dual gauge parameter of $\eta(G)$ given by:
$\eta^{\circ}(G):=\min\{\mu:\mu\geq 0,N\succcurlyeq 0,\text{diag}(N)=\mu\cdot\mathbbm{1},(1/4)\cdot L^{*}(N)\geq\mathbbm{1}\},$ (4)
where $L^{*}$ is a function that maps positive semidefinite matrices to a vector in +$\mathbb{R}^{E}_{+}$ such that $L^{*}(N)_{ij}=N_{ii}+N_{jj}-2N_{ij}$, which can be seen as the adjoint of the Laplacian function that maps nonnegative edge weights to the usual weighted Laplacian matrix. Using the framework of gauge duality, one can show that the pairs $\text{mc}(G),\text{fcc}(G)$ and $\eta(G),\eta^{\circ}(G)$ satisfy
$|E|$ $\leq\text{mc}(G)\text{fcc}(G),$ (5a)
$|E|$ $\leq\eta(G)\eta^{\circ}(G),$ (5b)
and it can also be shown that equality is attained in both cases for edge-transitive graphs (for a complete discussion of gauge duality and its applications to graph parameters, see [14], and for a study of the equality cases of similar equations as the above involving the Lovász Theta, see [8]).
In this context, we are interested in studying the optimal solutions for the semidefinite programs $\eta(G),\eta^{\circ}(G)$ when $G$ presents high structural regularity in order to better understand their relationship with maxcut and fcc. We are also interested in studying the solutions of more general semidefinite programs that can be used to approximate problems such as max 2-sat. More specifically, this work achieves the following:
  • In Sections 3.1, 3.2, we extend the equality cases of (5b) for graphs belonging to certain coherent configurations, which include all distance-regular graphs. We combine an observation from [2] about orthogonal projections onto *-algebras with the algebraic properties of coherent algebras to construct the proof;
  • In Section 3.3, we calculate $\eta^{\circ}(G)$ in terms of the smallest eigenvalue of the adjacency matrix of $G$ for graphs belonging to symmetric association schemes, which in turn yields a spectral bound to fcc. The proof utilizes a key observation from [11] that the structure of such graphs allows us to convert certain semidefinite programs into linear programs, enabling us to compute an explicit optimal solution;
  • In Section 4, we consider a more general semidefinite program framework used for approximating quadratic programs, first introduced in [12] to approximate max 2-sat. We compute the optimal value of this semidefinite program when the underlying graphs belong to a symmetric association scheme, which in turn yields bounds to max 2-sat in term of the first eigenmatrix of the scheme, and we also compute the optimum of its gauge dual when the scheme arises from a distance-regular graph. We combine the methods developed throughout Section 3 with classical facts about association schemes, such as the orthogonality relations and the algebraic properties of the eigenmatrices, to arrive at the desired results.
2 Coherent configurations and association schemes
We say that a set $\mathcal{C}=\{A_{0},A_{1},...,A_{d}\}$ of square matrices with entries in $\{0,1\}$ is a coherent configuration if the following hold:
  1. (1)
    $\sum_{i}A_{i}=J$, where $J$ denotes the matrix with all entries equal to one;
  2. (2)
    If $A_{i}\in\mathcal{C}$, then $A_{i}^{T}\in\mathcal{C}$;
  3. (3)
    If $A_{i}\in\mathcal{C}$ and $A_{i}$ has a nonzero diagonal entry, then it is a diagonal matrix;
  4. (4)
    There are nonnegative integers $p_{ij}^{l}$ such that
    $A_{i}A_{j}=\sum_{l=0}^{d}p_{ij}^{l}\cdot A_{l},$
    for any indices $0\leq i,j\leq d$.
The diagonal matrices in $\mathcal{C}$ that partition the identity matrix $I$ are called fibers, and if there is only one of these, that is, if $I\in\mathcal{C}$, then we say that the configuration is homogeneous [^1] Some authors will call these (non-commutative) association schemes [4], however we follow Cameron [7] in reserving the name association scheme to when all matrices in the configuration are symmetric. Note that an association scheme is necessarily a homogeneous coherent configuration.. Coherent configurations were initially introduced by Higman [13] to study permutation groups, as, for instance, one can always construct a coherent configuration from the orbitals of a permutation group acting on some set, as well as from the Cayley graphs generated by its conjugacy classes. Association schemes on the other hand were first introduced to study statistical experiments, but have since found many important applications in graph theory — e.g. in the study of distance-regular and strongly-regular graphs [6] —, coding theory — e.g. in the study of error-correcting codes [9] —, and in many other areas from combinatorics [3].
If the matrices in $\mathcal{C}$ commute, we say that the configuration is commutative. In this case, we get that the linear span of the matrices in $\mathcal{C}$ over $\mathbb{C}$ forms a commutative *-algebra — that is, an algebra closed under conjugate-transposition —, which in turn implies that there exists a basis of primitive idempotents $E_{0}=|V|^{-1}\cdot J,E_{1},...,E_{d}$ that are the projectors onto the common eigenspaces of the matrices in $\mathcal{C}$ (see for instance [3, Ch.2]). The *-algebra spanned by the matrices in a general configuration $\mathcal{C}$ is known as a coherent algebra, and in the case of association schemes we obtain the well-known Bose-Mesner algebra.
The familiar vertex-transitive and edge-transitive graphs are intimately related to coherent algebras, since if $\Gamma=\text{Aut}(G)$ denotes the automorphism group of a graph $X$, then the set
$C(\Gamma):=\{M\in M_{V}(\mathbb{C}):MP=PM,\forall P\in\Gamma\}$
known as the centralizer algebra (or commutant algebra) of $\Gamma$ is a coherent algebra, where the group $\Gamma$ is identified with its natural representation as a subgroup of permutation matrices. Some structural properties of $\Gamma$ correspond to algebraic properties of $C(\Gamma)$, for instance, if $G$ is vertex-transitive then $C(\Gamma)$ is always homogeneous, and if $G$ is edge-transitive we have three possibilities: (i) if $G$ is $1$-arc-transitive, then $C(\Gamma)$ will be homogeneous and $A$ will belong to $\mathcal{C}$; (ii) if $G$ is edge-transitive, vertex-transitive but not $1$-arc-transitive, then $C(\Gamma)$ will be homogeneous and $A$ will split in $\mathcal{C}$, that is, $A=A_{i}+A_{i}^{T}$ for some index $0\leq i\leq d$; or (iii) if $G$ is edge-transitive, but not vertex-transitive, then it will be bipartite and thus $C(\Gamma)$ will have two fibers and $A$ will split in $\mathcal{C}$ (see [5, Ch.17] for more details).
Our analysis relies on one key result about these objects: the orthogonal projection onto a coherent algebra preserves positive semidefiniteness, that is, if $M$ is positive semidefinite, then the orthogonal projection $M^{\prime}$ onto a coherent algebra $\mathcal{A}$ is also positive semidefinite (see [2, Corollary 9.1]). In the next sections, we use this fact to study solutions to (2) and (4) that lie in coherent algebras with similar structure as to the ones associated with edge-transitive graphs, that is, we consider graphs whose adjacency matrix $A$ either belongs to or splits in some coherent configuration $\mathcal{C}$.
3 MAXCUT and FCC
3.1 Graphs in coherent configurations
In this section we study the equality cases of (5b), which will prove to be useful for deriving the desired spectral bounds for fcc and max 2-sat. We begin with the following lemma:
Lemma 1.
Let $G$ be a graph with adjacency matrix $A$, and let $\mathcal{C}$ be a coherent configuration such that $A$ either belongs to or splits in $\mathcal{C}$. If $M$ is a feasible solution for (2), then its orthogonal projection onto the coherent algebra $\mathcal{A}$ generated by $\mathcal{C}$ is also feasible and has the same objective value. The same is true for any feasible solution of (4).
Proof.
By [2, Corollary 9.1], it follows that if $M$ is positive semidefinite then so is $M^{\prime}$, thus it suffices to show that $\text{diag}(M^{\prime})=\text{diag}(M)=\mathbbm{1}$. Let $\mathcal{C}=\{A_{0},...,A_{d}\}$, and since these matrices form an orthogonal basis to $\mathcal{A}$, we have that
$M^{\prime}=\sum_{i=0}^{d}\frac{\langle M,A_{i}\rangle}{\langle A_{i},A_{i}\rangle}\cdot A_{i}.$
If $A_{i}$ is a fiber of the configuration, hence a 01-diagonal matrix, we get that $\langle M,A_{i}\rangle=\langle A_{i},A_{i}\rangle$, implying that the diagonal entries of $M^{\prime}$ are all equal to one.
Now to see that the same holds for (4), it suffices to show that if $N$ is a feasible solution then $L^{*}(N^{\prime})\geq 4\cdot\mathbbm{1}$. Indeed, note that if $A_{i}\in\mathcal{C}$, then $\langle A_{i},N\rangle=\langle A_{i}^{T},N\rangle$ since $N$ is symmetric, and as $A$ either belongs to or splits in $\mathcal{C}$, it follows that $N^{\prime}_{ij}=N^{\prime}_{ji}=\langle A,N\rangle/\langle A,A\rangle$ for any edge $ij\in E$, that is, the orthogonal projection is constant on the edges of $G$. This in turn means that for any edge $ij\in E$ we have
$L^{*}(N^{\prime})_{ij}=2\mu-\frac{2\langle A,N\rangle}{\langle A,A\rangle}=2\mu-\frac{4\sum_{uv\in E}N_{uv}}{2m},$
and then by combining this with the fact that $L^{*}(N)_{ij}=2\mu-2N_{ij}\geq 4$, we get that $L^{*}(N^{\prime})_{ij}\geq 4$ for any edge $ij\in E$, as desired. ∎
The previous result implies that we may always assume that optimal solutions for (2) and (4) lie in $\mathcal{A}$. We can also conclude a type of equivalence between the feasibility conditions of the solutions of these programs.
Lemma 2.
Let $G$ be a graph as in Lemma 1, and let $M\in\mathcal{A}$. Then for any $\mu>0$, $N=\mu\cdot M$ is a feasible solution for (4) if, and only if, $M$ is a feasible solution for (2) with objective value at least $|E|/\mu$.
Proof.
The key point here is to show the equivalence between the constraint $L^{*}(N)\geq 4\cdot\mathbbm{1}$ and $\langle L,N\rangle\geq 4|E|$. Indeed, note that if $N$ belongs to $\mathcal{A}$, then by the observations made in the proof of the previous lemma it follows that the entries of $N$ are constant on the edges of $G$, thus $L^{*}_{ij}\geq 4$ for any $ij$ if and only if $\sum_{ij\in E(G)}L^{*}_{ij}(N)\geq 4|E|$. Noting that we can view $L$ as the weighted Laplacian function evaluated at $\mathbbm{1}$ with adjoint given by $L^{*}$, we get that this constraint is equivalent to
$\langle L,N\rangle=\langle\mathbbm{1},L^{*}(N)\rangle=\sum_{ij\in E(G)}L^{*}_{ij}(N)\geq 4|E|,$
and with this the result immediately follows. ∎
Combining the two lemmas, we prove the main result of this section:
Theorem 3.
If $G$ is a graph whose adjacency matrix either belongs to or splits in a coherent configuration $\mathcal{C}$, then
$\eta(G)\eta^{\circ}(G)=|E|.$
Proof.
Let $M$ be an optimal solution for (2), which by Lemma 1 can be assumed to belong to the coherent algebra generated by $\mathcal{C}$. We can thus take $\mu=|E|/\eta\geq 0$ and consider the matrix $N=\mu\cdot M$, which is a feasible solution for the gauge dual in (4) by Lemma 2 and the fact that $\langle L/4,M\rangle=|E|/\mu$. Also, $N$ has objective value $|E|/\eta$ for (4), and as this is a minimization problem, we get that
$\eta(G)\eta^{\circ}(G)\leq|E|,$
which together with (5b) implies the desired result. ∎
We recall that if $G$ is edge-transitive, and if $\Gamma$ is its automorphism group, then the centralizer $C(\Gamma)$ is a coherent algebra whose coherent configuration either contains or splits $A$. In the case of distance-regular graphs, their Bose-Mesner algebra clearly contains $A$ in its association scheme. Thus, we obtain the following corollary:
Corollary 4.
If $G$ is either an edge-transitive or a distance-regular graph, then (5b) is an equality.
$\blacksquare$
3.2 1-walk-regular graphs
We say that a graph $G$ is $1$-walk-regular if, for any nonnegative integer $k$, there exists constants $a_{k},b_{k}$ such that
$A^{k}\circ I=a_{k}\cdot I\quad\text{and}\quad A^{k}\circ A=b_{k}\cdot A.$
This definition generalizes graphs that are both vertex-transitive and edge-transitive, but also includes for instance all distance-regular graphs. We study this family of graphs individually since their adjacency matrices do not belong to nor split in coherent configurations in general, but they still have enough structure for us to prove the desired result. We may consider the algebra generated by the adjacency matrix of such graphs, note that it is a *-algebra, and we can find an orthogonal basis $\{I,A,A_{2},...,A_{d}\}$ such that for $k\geq 2$ we have $A_{k}\circ I=0$ and $A_{k}\circ A=0$. All matrices of $\mathcal{A}$ have constant diagonal entries and are also constant on the edges of $G$. This means we can also assume that any optimal solution $M$ to (2) lies in this algebra, and moreover its entries $M_{ij}$ corresponding to any edge $ij\in E$ will be given by $M_{ij}=M_{ji}=\langle M,A\rangle/\langle A,A\rangle$, hence the same reasoning applied in the proof of Theorem 3 allows us to conclude the following:
Corollary 5.
If $G$ is a $1$-walk-regular graph, then (5b) is an equality.
$\blacksquare$
3.3 Association schemes
We are now interested in studying the semidefinite programs (2) and (4) for graphs belonging to association schemes. Throughout this section, we let $G$ be a $k$-regular graph with adjacency matrix $A$ belonging to an association scheme $\mathcal{S}=\{I,A,A_{2},...,A_{d}\}$ with Bose-Mesner algebra $\mathcal{A}$ and primitive idempotents $E_{0},...,E_{d}$, and we let $P,Q\in M_{d+1}(\mathbb{R})$ be the first and second eigenmatrices of $\mathcal{S}$, respectively, that is, the matrices defined by the equations
$A_{i}=\sum_{l=0}^{d}P_{li}\cdot E_{l}\quad\text{and}\quad E_{i}=\frac{1}{|V|}\sum_{l=0}^{d}Q_{li}\cdot A_{l}.$ (6)
We note that $PQ=QP=|V|\cdot I$, and that these are both real matrices as $\mathcal{S}$ is a set of symmetric matrices. We denote the degrees of the scheme by $k_{i}=p_{ii}^{0}$, hence each matrix $A_{i}$ can be seen as the adjacency matrix of a $k_{i}$-regular graph, and we further note that $P_{l0}=Q_{l0}=1$, $P_{0l}=k_{l},Q_{0l}=m_{l}$,$P_{ji}m_{j}=Q_{ij}k_{i}$, for all indices $0\leq i,j,l\leq d$, where $m_{l}=\text{tr}(E_{l})$ (see [3, Ch.2] for more details). From these definitions one can obtain the well-known orthogonality relations:
$\sum_{l=0}^{d}\frac{P_{il}P_{jl}}{k_{l}}=\begin{cases}\frac{|V|}{m_{i}}&\quad\text{if }i=j,\\ 0&\quad\text{otherwise},\end{cases}\quad\text{and}\quad\sum_{l=0}^{d}P_{li}P_{lj}m_{l}=\begin{cases}|V|k_{i}&\quad\text{if }i=j,\\ 0&\quad\text{otherwise}.\end{cases}$
From (6), the eigenvalues of any linear combination $\sum_{i}x_{i}\cdot A_{i}$ of the basis matrices of $\mathcal{A}$ are given by the linear combinations of the respective eigenvalues of the $A_{i}$, hence the constraint $M\in\mathcal{A},M\succcurlyeq 0$ is equivalent to requiring that $Px\geq 0$, where $x\in\mathbb{R}^{d+1}$ is the vector of coefficients such that $M=\sum_{i}x_{i}\cdot A_{i}$. With these observations, we can prove the following:
Theorem 6.
If $G$ is a $k$-regular graph and whose adjacency matrix $A$ belongs to an association scheme, and if $\lambda_{\text{min}}(A)$ is its smallest eigenvalue, then
$\eta^{\circ}(G)=\frac{2k}{k-\lambda_{\text{min}}(A)}.$
In particular, we have
$\frac{2k}{k-\lambda_{\text{min}}(A)}\leq\text{fcc}(G)\leq\frac{1}{\alpha_{GW}}\left(\frac{2k}{k-\lambda_{\text{min}}(A)}\right).$
Proof.
The previous comments about association schemes combined with Lemma 1 allows us to rewrite the semidefinite program $\eta^{\circ}(G)$ as the linear program
$\eta^{\circ}(G)=\min\left\{x_{0}:x\in\mathbb{R}^{d+1},Px\geq 0,x_{0}\geq 0,x_{0}-x_{1}\geq 2\right\},$
where $x=(x_{0},...,x_{d})$, and the constraint $x_{0}-x_{1}\geq 2$ follows from the constraint $L^{*}(N)\geq 4\mathbbm{1}$ and the fact that $L^{*}(N)_{ij}=2x_{0}-2x_{1}$ if we assume that $N\in\mathcal{A}$. We can then use linear program strong duality and some straightforward algebraic manipulation to obtain:
$\eta^{\circ}(G)=\max\left\{2b:y\in\mathbb{R}_{+}^{d+1},a,b\in\mathbb{R}_{+},P^{T}y=b\cdot e_{1}+(1-a-b)\cdot e_{0}\right\},$
where $e_{i}$ is the $i$-th canonical basis vector in $\mathbb{R}^{d+1}$, with indexing starting at 00. Noting that the inverse of $P$ is $(1/|V|)\cdot Q$, that $P_{ji}m_{j}=Q_{ij}k_{i}$ and that the first row of $Q$ contains the traces of the respective primitive idempotents, we get
$y=\frac{1}{|V|}\sum_{l=0}^{d}\left(\left(\frac{bP_{l1}}{k}+1-a-b\right)m_{l}\right)\cdot e_{l},$
thus $y$ is completely determined by our choice of $a,b$. We can also note that if $(y,a,b)$ is feasible for our program, we can take $a^{\prime}=0$ and $y^{\prime}=y+(a/|V|)\cdot Q^{T}e_{0}$, which is nonnegative since $Q_{0i}=m_{i}\geq 0$ for any index $0\leq i\leq d$, and thus $(y^{\prime},a^{\prime}=0,b)$ is a feasible solution with the same objective value. So we may always assume that an optimal solution has $a=0$. Combining this with the fact that $P_{01}=k$ and that each entry of $y$ must be nonnegative, we get that
$b\leq\frac{k}{k-P_{l1}},$
for all indices $1\leq l\leq d$. As we wish to maximize $b$, we set
$b=\underset{1\leq l\leq d}{\min}\left\{\frac{k}{k-P_{l1}}\right\}=\frac{k}{k-\lambda_{\text{min}}(A)},$
and with this the result follows immediately. ∎
We remark that one could also obtain the previous result by combining Theorem 3 with the result from [11] that $\eta(G)=(|V|/4)(k-\lambda_{\text{min}}(A))$ for graphs in association schemes. However, the constructive proof given above highlights some methods that will be particularly useful in the next section.
4 Quadratic programs and MAX 2-SAT
We now study a broader framework for approximating quadratic programs via semidefinite programs, that generalizes our previous work for maxcut. We let $G_{1}=(V,E_{1}),G_{2}=(V,E_{2})$ be graphs on the same vertex set, and set $L$ to be the Laplacian of $G_{1}$, and $K$ to be the signless laplacian of $G_{2}$. From this, we define the following quadratic program:
$\text{qp}(G_{1},G_{2}):=\max\left\{\left\langle\frac{L+K}{2},xx^{T}\right\rangle:x\in\mathbb{R}^{V},x_{i}^{2}=1\right\}.$ (7)
If we consider $G_{2}$ as the empty graph, then $\text{mc}(G_{1})=(1/2)\text{qp}(G_{1},G_{2})$, and we can similarly consider the semidefinite relaxation of this program given by
$\gamma(G_{1},G_{2}):=\max\left\{\left\langle\frac{L+K}{2},M\right\rangle:M\succcurlyeq 0,\text{diag}(M)=\mathbbm{1}\right\},$ (8)
where $\eta(G_{1})=(1/2)\gamma(G_{1},G_{2})$ in case $G_{2}$ is the empty graph. In [12], Goemans and Williamson showed how to adapt their maxcut algorithm in order to obtain an $\alpha_{\text{GW}}$-approximation for $\text{qp}(G_{1},G_{2})$ using an optimal solution of $\gamma(G_{1},G_{2})$, and then showed how to model max 2-sat with $\text{qp}(G_{1},G_{2})$. Similarly to (4), we define the program
$\gamma^{\circ}(G_{1},G_{2}):=\min\{\mu:$ $\mu\geq 0,N\succcurlyeq 0,\text{diag}(N)=\mu\cdot\mathbbm{1},$
$(1/2)\cdot L^{*}(N)\geq\mathbbm{1},(1/2)\cdot K^{*}(N)\geq\mathbbm{1}\},$ (9)
where $K^{*}$ is a function that maps a positive semidefinite matrix $N$ to a vector in $\mathbb{R}_{+}^{E_{2}}$ such that $K^{*}(N)_{ij}=N_{ii}+N_{jj}+2N_{ij}$ — that is, the adjoint of the signless Laplacian function that maps nonnegative edge weights to the usual weighted signless Laplacian matrix. The parameters $\gamma,\gamma^{\circ}$ form a primal-dual pair of gauges (when specializing to the unweighted scenario, as discussed in the introduction) satisfying the inequality
$|E_{1}|+|E_{2}|\leq\gamma(G_{1},G_{2})\gamma^{\circ}(G_{1},G_{2}),$ (10)
and similarly to what was done in the previous sections, we inquiry into the optimal solutions of these semidefinite programs for graphs with high structural regularity. We also discuss the equality cases of (10).
From now on, we assume that the adjacency matrices $A_{1},A_{2}$ of $G_{1},G_{2}$ are elements of an association scheme $S=\{I,A_{1},A_{2},...,A_{d}\}$ with Bose-Mesner algebra $\mathcal{A}$, and we use the same notation as in Section 3.3 to refer to the primitive idempotents and eigenmatrices of $\mathcal{S}$. We first note that the feasibility region of (8) is the same as (2), hence any feasible solution $M$ may be assumed to belong to $\mathcal{A}$ by an analogous version of Lemma 1. We can apply a similar argument as in Section 3.3 to transform $\gamma(G_{1},G_{2})$ into a linear program: the constraint $\text{Diag}(M)=\mathbbm{1}$ implies that if $M\in\mathcal{A}$ then $M=I+\sum_{l=1}^{d}x_{i}\cdot A_{i}$, hence the constraint $M\succcurlyeq 0$ is equivalent to $Rx\geq 0$, where $x\in\mathbb{R}^{d}$ and $R$ is the $(d+1)\times d$ matrix obtained by deleting the first column of $P$. Noting that $L=k_{1}\cdot I-A_{1},Q=k_{2}\cdot I+A_{2}$, it follows that
$\frac{L+Q}{2}=\frac{1}{2}(\langle k_{1}\cdot I-A_{1},M\rangle+\langle k_{2}\cdot I+A_{2},M\rangle)=\frac{|V|}{2}((k_{1}+k_{2})+(k_{2}x_{2}-k_{1}x_{1})),$
and with this we can rewrite (8) as the following linear program:
$\begin{split}\gamma(G_{1},G_{2})&=\frac{|V|}{2}((k_{1}+k_{2})+\max\{(k_{2}\cdot e_{2}-k_{1}\cdot e_{1})^{T}x:x\in\mathbb{R}^{d},Rx\geq-\mathbbm{1}\})\\ &=\frac{|V|}{2}((k_{1}+k_{2})+\min\{\mathbbm{1}^{T}y:y\in\mathbb{R}^{d+1}_{+},R^{T}y=k_{1}\cdot e_{1}-k_{2}\cdot e_{2}\}),\end{split}$ (11)
where the second equality follows from strong duality. As a consequence, we explicitly compute $\gamma(G_{1},G_{2})$:
Theorem 7.
If $G_{1},G_{2}$ are graphs whose adjacency matrices $A_{1},A_{2}$ belong to an association scheme with first eigenmatrix $P$, then
$\gamma(G_{1},G_{2})=\frac{|V|}{2}\left((k_{1}+k_{2})+\underset{0\leq l\leq d}{\max}\{P_{l2}-P_{l1}\}\right).$
Proof.
We let $z$ denote the first row of the second eigenmatrix $Q$, that is, we have that ${z=(1,m_{1},...,m_{d})}$ is a vector containing the multiplicities of the scheme, and moreover, since $QP=|V|\cdot I$, and $R$ is obtained from deleting the first column of $P$, it follows that $R^{T}z=0$. As $P$ is invertible, we get that $R^{T}$ has rank $d$, which implies that $\ker(R^{T})=\langle z\rangle$. On the other hand, we may use the orthogonality relations to explicitly construct a solution for the system $R^{T}y=k_{1}\cdot e_{1}-k_{2}\cdot e_{2}$ by considering
$y^{\prime}=\frac{1}{|V|}\sum_{l=0}^{d}((P_{l1}-P_{l2})m_{l})\cdot e_{l},$
and then noting that
$(R^{T}y^{\prime})_{i}=\frac{1}{|V|}\left(\sum_{l=0}^{d}P_{l1}P_{li}m_{l}-\sum_{l=0}^{d}P_{l2}P_{li}m_{l}\right)=\begin{cases}k_{1},&\quad\text{if }i=1,\\ -k_{2},&\quad\text{if }i=2,\\ 0,&\quad\text{otherwise}.\\ \end{cases}$
This implies that any solution for $R^{T}y=k_{1}\cdot e_{1}-k_{2}\cdot e_{2}$ is of the form $y=y^{\prime}+(\alpha/|V|)\cdot z$, for any real $\alpha$, but note also that $\mathbbm{1}^{T}y^{\prime}=0$ by the orthogonality relations, and $\mathbbm{1}^{T}z=\sum_{l=0}^{d}m_{l}=|V|$. Hence the objective value of any feasible $y$ is precisely $\alpha$. We then need to choose the minimum $\alpha$ that makes $y$ feasible, which is given by
$\alpha^{*}=\underset{0\leq l\leq d}{\max}\{P_{l2}-P_{l1}\},$
since the entries of $y$ must all be nonnegative, and this together with (11) concludes the proof. ∎
The previous proof combined with the approximation algorithm given in [14] gives bounds to max 2-sat in terms of the eigenvalues of $G_{1}$ and $G_{2}$ when they belong to an association scheme.
We turn our attention to $\gamma^{\circ}(G_{1},G_{2})$, and in order to compute its optimal value analytically, we shall restrict ourselves to association schemes coming from distance-regular graphs. First, we recall that if a distance-regular graph $G$ with diameter $d$ has intersection array $\iota(G)=\{b_{0},...,b_{d-1};c_{1},...,c_{d}\}$, and if $P$ is the first eigenmatrix of the scheme generated by $G$, then
$P_{l2}=\frac{k_{2}}{b_{1}k_{1}}(P_{l1}^{2}-(k_{1}-b_{1}-1)P_{l1}-k_{1}),$ (12)
for any index $0\leq l\leq d$ (see [6, Ch.4] for more details). With this, we prove the following:
Theorem 8.
Let $G_{1}$ be a distance-regular graph with diameter $d$ and let $G_{2}$ be its distance-$2$ graph, with respective adjacency matrices $A_{1}$ and $A_{2}$. If $P$ is the first eigenmatrix associated with the symmetric scheme generated by $A_{1}$, then
$\gamma^{\circ}(G_{1},G_{2})=\begin{cases}\frac{k_{1}}{k_{1}-P_{d1}},&\quad\text{if }k_{2}P_{d1}+k_{1}P_{d2}>0,\\ \frac{k_{1}}{k_{1}-P_{d1}}-\frac{(k_{2}P_{d1}+k_{1}P_{d2})}{2k_{2}(k_{1}-P_{d1})},&\quad\text{otherwise.}\end{cases}$
Proof.
Similarly to what was done in Theorem 6, we can rewrite $\gamma^{\circ}(G_{1},G_{2})$ as the following linear program:
$\gamma^{\circ}(G_{1},G_{2})=\min\{x_{0}:x\in\mathbb{R}^{d+1},Px\geq 0,x_{0}-x_{1}\geq 1,x_{0}+x_{2}\geq 1\},$
where the last two constraints follow from the fact that any feasible $N\in\mathcal{A}$ for (4) is constant on the edges of $G_{1}$ and $G_{2}$. By strong duality, we have
$\gamma^{\circ}(G_{1},G_{2})=\max\{b+c:y\in\mathbb{R}^{d+1}_{+},a,b,c\in\mathbb{R}_{+},P^{T}y=b\cdot e_{1}-c\cdot e_{2}+(1-a-b-c)\cdot e_{0}\},$ (13)
and since $|V|^{-1}\cdot Q$ is the inverse of $P$ and $P_{ji}m_{j}=Q_{ij}k_{i}$, we get that
$y=\frac{1}{|V|}\sum_{l=0}^{d}\left(\left(\frac{bP_{l1}}{k_{1}}-\frac{cP_{l2}}{k_{2}}+1-a-b-c\right)m_{l}\right)\cdot e_{l},$
and so $y$ is determined by our choice of $a,b,c$. Similarly to the proof of Theorem 6, we may always assume that any optimal solution $(y,a,b,c)$ is such that $a=0$, and since each entry of $y$ must be nonnegative, we get that $20\leq c\leq 1/$ 2, and that
$b\leq\underset{1\leq l\leq d}{\min}\left\{\left(\frac{k_{1}}{k_{1}-P_{l1}}\right)\left(\frac{(1-c)k_{2}-cP_{l2}}{k_{2}}\right)\right\}.$
If we fix some feasible $c$, as we wish to maximize $b+c$, we may always assume that the previous inequality is an equality, and thus any solution to our program is in fact uniquely determined by our choice of $c$. A simple algebraic manipulation then shows that we can rewrite $\gamma^{\circ}(G_{1},G_{2})$ as
$\gamma^{\circ}(G_{1},G_{2})=\underset{0\leq c\leq 1/2}{\max}\left\{\underset{1\leq l\leq d}{\min}\left\{\frac{k_{1}}{k_{1}-P_{l1}}-c\frac{(k_{1}P_{l2}+k_{2}P_{l1})}{(k_{1}-P_{l1})k_{2}}\right\}\right\}.$ (14)
We now claim that for any fixed $c$, the index $l$ that achieves the minimum in the previous equation is $d$, and we shall prove this by constructing a chain of equivalent inequalities. Indeed, for any index $1\leq l\leq d$, define
$\gamma^{\circ}_{l}(c)=\frac{k_{1}}{k_{1}-P_{l1}}-c\frac{(k_{1}P_{l2}+k_{2}P_{l1})}{(k_{1}-P_{l1})k_{2}},$
and assume w.l.o.g. that the matrix $P$ is ordered such that $P_{d1}<\ldots<P_{01}$, that is, we order the distinct eigenvalues of $G_{1}$ in decreasing order (and note that there are precisely $d+1$ of those as $G_{1}$ is distance-regular with diameter $d$). Now consider the indices $1\leq j\leq i\leq d$, and note that it suffices to show that $\gamma^{\circ}_{i}(c)\leq\gamma^{\circ}_{j}(c)$ for a fixed $20\leq c\leq 1/$ 2. A straightforward computation shows that this is equivalent to requiring that
$P_{j2}(k_{1}-P_{i1})-P_{i2}(k_{1}-P_{j1})\leq k_{2}(P_{j1}-P_{i1}).$
By substituting the terms $P_{j2}$ and $P_{i2}$ with the respective expressions given in (12) and noting that by assumption $P_{j1}-P_{i1}>0$, we get that the previous inequality is equivalent to $P_{i1}(k_{1}-P_{j1})\leq k_{1}(k_{1}-P_{j1})$, which holds since $k_{1}$ is the largest eigenvalue of $G_{1}$ and $j>0$. By combining this fact with (14), we obtain
$\gamma^{\circ}(G_{1},G_{2})=\underset{0\leq c\leq 1/2}{\max}\left\{\frac{k_{1}}{k_{1}-P_{d1}}-c\left(\frac{k_{2}P_{d1}+k_{1}P_{d2}}{k_{2}(k_{1}-P_{d1}}\right)\right\}.$
If $k_{2}P_{d1}+k_{1}P_{d2}>0$, we take $c=0$ in order to maximize the objective value, and otherwise we take $c=1/2$, which concludes the proof. ∎
We remark that one could further reduce the expression obtained in the previous theorem for $\gamma^{\circ}$ by substituting $P_{d2}$ according to (12), which would result in a term solely dependent on $P_{d1}$. This however results in a rather obscure expression that does not seem to present any further insight into the problem, and so we opted to omit it.
Since the semidefinite programs for $\gamma$ and $\gamma^{\circ}$ generalize $\eta$ and $\eta^{\circ}$, a natural question to ask is whether the equality cases of (10) behave similarly to the ones of (5b). It turns out that even when we take $G_{1}$ to be distance-regular and edge-transitive, and $G_{2}$ to be its distance-2 graph, there are examples where (10) is a strict inequality. Take for instance the Paley graph of parameter $9$, obtained by first considering a finite field $\mathbb{F}_{9}$ with $9$ elements which will be the vertices of our graph, and then connecting two vertices if their difference is a square in $\mathbb{F}_{9}$. This is an arc-transitive graph and also a strongly regular graph, hence its adjacency matrix $A$ belongs a symmetric association scheme $\{I,A,J-A-I\}$, with first eigenmatrix given by
$P=\begin{pmatrix}1&4&4\\ 1&1&-2\\ 1&-2&1\end{pmatrix}.$
If we then let $G_{1},G_{2}$ be such graph and its complement, respectively, one can easily see that the vector $y=(0,1/2,0,0,1/2,1/4)$ is a feasible solution for the maximization program in (13) with objective value $3/4>8/11=(|E_{1}|+|E_{2}|)/\gamma$, implying that $\gamma(G_{1},G_{2})\gamma^{\circ}(G_{1},G_{2})>|E_{1}|+|E_{2}|$.
5 Conclusion
In this work, we discussed semidefinite programs used to approximate certain NP-hard graph parameters, and showed that they can be explicitly computed in terms of the eigenvalues when the underlying graphs present enough structural regularity. These computations yielded spectral bounds for fcc and max 2-sat, and also provided new insights into the relationship between these semidefinite programs and their combinatorial counterparts.
Here are some possible future research directions. We believe that the results from Sections 3.1 and 3.2 can provide further insight into the equality cases of (5a), which in turn could lead to a deeper understanding of the relationship between maxcut and fcc. The example discussed in Section 4 suggests that stronger conditions on the underlying graphs of max 2-sat may be required to attain equality in (10), and future inquiry into these conditions could provide new interesting families of graphs and association schemes. Determining the combinatorial structure of the gauge dual of max 2-sat may also lead to interesting results, and investigating whether one can construct an approximation algorithm for this parameter with $\gamma^{\circ}$ could lead to spectral bounds by means of the results in the previous section.
Acknowledgements
Both authors acknowledge the support from FAPEMIG and CNPq.